### v1

# Universidad Andres Bello, Facultad de Ciencias Exactas, Departamento de Ciencias Físicas - Instituto de Astrofísica, Autopista Concepcion-Talcahuano 7100, Talcahuano, Chile Instituto de Física, Universidad de Antioquia, Calle 70 52-21, Medellín, Colombia

Abstract
Context. The Bulge Globular Cluster NGC 6316 has been the subject of a limited amount of previous chemical studies besides metallicity. As part of the bulge Cluster APOgee Survey (CAPOS), we can now improve our knowledge of the chemistry and nature of this cluster. CAPOS makes use of high-resolution, high Signal-to-Noise Ratio Near-Infrared spectroscopy, which is capable of penetrating the substantial dust towards the Galactic Bulge - a significant obstacle for this cluster due to its high reddening.
Aims. We aim to conduct the most robust chemical study to date for NGC 6316. Using high-resolution, high Signal-to-Noise Ratio spectra of 10 RGB members, we aim to derive abundances for a number of elements with a variety of nucleosynthetic origins, most of which have never been studied before in this cluster.
Methods. We use the Brussels Automatic Code for Characterizing High accuracy Spectra (BACCHUS) with atmospheric parameters photometrically obtained in order to determine the abundances.
Results. We determined, for the first time, high resolution spectroscopic abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce for this cluster. We obtained a mean metallicity [Fe/H] = 0.87 0.02, finding no indication of an intrinsic metallicity spread. Our metallicity agrees with the most recent values from other studies, revising earlier values that were 0.5 dex metal-richer. With this new value, this cluster, long believed to be a member of the classical metal-rich group of bulge GCs around -0.5, now falls in the dominant bulge globular cluster peak around [Fe/H] = -1. The cluster presents a clear C-N anticorrelation. We also found a [/Fe] = 0.3 0.02. Our abundances show similar behaviour to other in situ globular clusters with comparable metallicity. We obtained and by isochrone fitting, in good agreement with the recent determinations from other works. We derive an overall metallicity by isochrone fitting, in agreement with our abundance determination. According to the mean [Mg/Fe] and [Al/Fe] abundances from first population stars, NGC 6316 is an in-situ globular cluster, in accordance with various dynamical classifications.
Conclusions.
Key Words.:
Globular Clusters – Spectroscopy – Abundances1 Introduction
Globular Clusters (GCs) are among the oldest objects in the universe and also present Multiple Populations (MPs). These characteristics make GCs very interesting objects to study in the context of galactic and cluster formation. While studies related to the age of GCs have existed for quite some time now (Sandage, 1953), studies of MPs are relatively new. This concept, described as significant intrinsic abundance variations within a GC, emerged some 15 years ago when Carretta et al. (2009) suggested that this phenomenon must be fundamentally related to the mechanisms of formation and early evolution of GCs since, at that time, all the GCs studied (e.g., Bedin et al. (2004), among others) presented MPs. To this day, many GCs have been studied by combining different photometric (Frelijj et al., 2017, 2022; Piotto et al., 2015; Milone et al., 2017) and spectroscopic (Carretta et al., 2009; Mészáros et al., 2020) techniques. Great advances have been made, such as the creation of large databases (e.g., Baumgardt111https://people.smp.uq.edu.au/HolgerBaumgardt/globular/ and Harris, 1996 Galactic GCs databases), identification of chemical correlations and anti-correlations, differentiation of the populations through them (Carretta et al. (2010), Mészáros et al. (2020)), and even the identification of a new type of GC, Type II (Milone et al., 2017). However, the lack of a viable formation scenario for GCs and MPs, e.g. to account for the mass budget problem, as well as an explanation for the concomitant existence of at least some true Simple Stellar Population GCs (E3: Salinas, & Strader (2015); Monaco et al. (2018), Ruprecht 106: Villanova et al. (2013); Dotter et al. (2018); Frelijj et al. (2021), Ter7 & Pal12: Cohen (2004); Tautvaišienė et al. (2004); Sbordone et al. (2007)), among other issues, only proves that there is still a long way to go to fully understand the formation mechanisms of GCs. The only way to solve these questions is to obtain both more data, to help constrain formation scenarios, as well as additional theoretical ideas to test.
APOGEE ID | RA(J2000) | DEC(J2000) | G | J | H | Ks | pmRA | pmDEC | RV | SNR | ||
(deg) | (deg) | (masyr) | (masyr) | (kms) | (pixel) | |||||||
2M171638642809385 | 259.161017 | 15.309 | 16.458 | 14.246 | 12.549 | 11.582 | 11.329 | 94 | 65 | |||
2M171652352809502 | 259.218166 | 15.351 | 16.521 | 14.267 | 12.421 | 11.571 | 11.336 | 101 | 70 | |||
2M171636232808067 | 259.150986 | 14.713 | 15.835 | 13.478 | 11.55 | 10.213 | 10.212 | 99 | 80 | |||
2M171633302808396 | 259.138767 | 14.957 | 16.115 | 13.800 | 11.994 | 10.944 | 10.758 | 102 | 85 | |||
2M171640482808443 | 259.168675 | 15.210 | 16.262 | 14.045 | 12.274 | 11.325 | 11.046 | 103 | 85 | |||
2M171644822808302 | 259.186784 | 15.315 | 16.474 | 14.242 | 12.475 | 11.544 | 11.210 | 101 | 90 | |||
2M171633932811052* | 259.141393 | 14.742 | 15.988 | 13.624 | 11.496 | 10.646 | 10.375 | 103 | 105 | |||
2M171639032807212* | 259.162634 | 14.683 | 16.009 | 13.555 | 11.633 | 10.574 | 10.319 | 94 | 135 | |||
2M171639112804506 | 259.162961 | 14.549 | 15.962 | 13.357 | 11.282 | 10.239 | 9.875 | 102 | 140 | |||
2M171636272807166 | 259.151135 | 14.247 | 15.884 | 13.017 | 10.923 | 9.838 | 9.503 | 103 | 170 | |||
2M171626722811418 | 259.111349 | 16.483 | 17.576 | 15.454 | 13.796 | 12.909 | 12.763 | 94 | 30 | |||
2M171639922805581 | 259.166366 | 16.636 | 17.611 | 15.478 | 13.834 | 12.943 | 12.749 | 97 | 30 | |||
2M171650822807196 | 259.211757 | 16.280 | 17.245 | 15.278 | 13.660 | 12.855 | 12.573 | 98 | 35 | |||
2M171636222812074 | 259.150947 | 15.717 | 16.789 | 14.691 | 12.976 | 12.165 | 11.894 | 102 | 40 | |||
2M171641492807547 | 259.172901 | 15.539 | 16.584 | 14.478 | 12.839 | 11.959 | 11.648 | 92 | 45 | |||
2M171628822807520 | 259.120092 | 16.138 | 17.257 | 15.104 | 13.393 | 12.534 | 12.278 | 95 | 50 | |||
2M171644372806426 | 259.184902 | 16.061 | 17.137 | 15.050 | 13.406 | 12.567 | 12.324 | 107 | 55 | |||
2M171623772809447 | 259.099078 | 15.327 | 16.483 | 14.254 | 12.540 | 11.602 | 11.339 | 103 | 55 | |||
2M171635522806556 | 259.148019 | 15.473 | 16.573 | 14.422 | 12.753 | 11.818 | 11.609 | 105 | 55 |
The order of the columns is as follows: Star ID, Right Ascension(J2000), Declination(J2000), GAIA magnitude in G, GBP and GRP, 2MASS magnitude in J, H and Ks, Absolute Proper Motion in RA and DEC from GAIA, Heliocentric RV from ASPCAP, and SNR. IDs marked with an * indicate spectra not obtained by CAPOS.
Given that many of these questions remain unanswered after studying a large number of primarily halo GCs, a promising direction for future studies is to turn to the study of GCs in other Galactic components. The Galaxy possesses no less than three GC systems: the halo, bulge and disk systems (e.g. Massari, Koppelman, & Helmi, 2019). Indeed, these latter two contain fully 2/3 of the total number of Galactic GCs (Belokurov & Kravtsov, 2024). The study of bulge globular clusters (BGCs) in particular opens up the regime of high metallicity to MP studies since they include the highest metallicity of all GCs, and thus help verify metallicity trends so far only hinted at (e.g. Muñoz et al., 2020). Nevertheless, most of the Bulge/Disk (B/D) GC systems have not been studied in the detail required to quantify or even explore their MP nature, which demands either deep UV photometry with high spatial resolution, exemplified by the HST UV Legacy Survey of GCs Piotto et al. (2015), or high resolution high Signal-to-Noise Ratio (SNR) spectroscopy of large samples. Unfortunately, due to the vast amount of extinction in the plane and towards the center of the Galaxy (e.g Geisler et al., 2021), UV observations of the B/D GC systems are severely limited, so that the HST UV Legacy Survey includes only a very small number of these objects. Optical spectroscopy is similarly limited. However, the high reddening problem that previously prevented us from studying the B/D has been quite successfully addressed with the help of near-infrared (NIR) instruments. Surveys like the VISTA Variables in the Via Lactea (VVV) (Minniti et al., 2010) with its eXtension (VVVX)(Saito et al., 2024), and the Apache Point Observatory Galactic Evolution Experiment of the Sloan Digital Sky Survey III (SDSS-III APOGEE) (Majewski et al., 2017) with its successor (SDSS-IV APOGEE-2) capitalize on this capability to great success. However, despite these large surveys, the VVV survey lacks UV observations while the SDSS-IV survey did not prioritize B/D GCs. Mészáros et al. (2020) present the SDSS-IV sample of 44 GCs, of which only 2 are bonafide BGCs according to Massari, Koppelman, & Helmi (2019) and have a large enough sample of well-observed members and low enough reddening to qualify for their study. This is ¡4% of the total number of BGCs known, a disconcertingly low value.
In order to take full advantage of the wealth of astrophysical detail these key objects can provide, as complete a sample as possible is essential. To this end, ”the bulge Cluster APOgee Survey” (CAPOS) (Geisler et al., 2021) was created. This survey was designed to obtain detailed abundances and kinematics for as complete a sample of bona fide BGCs as possible using the unique advantages of APOGEE in order to exploit their extraordinary Galactic archaeology attributes. In addition to providing reliable multi-element abundances for a number of B/D GCs (many for the first time), this ongoing project has published a series of important results: e.g. confirmation of the scenario that Main Bulge and Main Disk GCs, formed in situ, have [Si/Fe] abundances slightly higher than their accreted counterparts at the same metallicity (Geisler et al., 2021, CAPOS I Paper); the existence of the characteristic N-C anticorrelation, and Al-N correlation for the first time in FSR 1758 (Romero-Colmenares et al., 2021, CAPOS II Paper); a Ce-N correlation in Tonantzintla 2 (Fernández-Trincado et al., 2022, CAPOS III Paper); and a N-C anticorrelation for the first time in NGC 6558 (González-Díaz et al., 2023, CAPOS IV Paper). These studies, which are providing valuable information about B/D GCs, demonstrate that it is possible to obtain and analyze excellent data that, until very recently, was very difficult to obtain.
Although there is already an APOGEE pipeline that derives chemical abundances and other parameters like radial velocity for BGCs, namely the APOGEE Stellar Parameters and Chemical Abundance Pipeline (ASPCAP, García Pérez et al. (2016)), it is known that the pipeline does not perform optimally for stars with extreme types of “non-standard” elemental abundance patterns, namely so-called Second Population (2P) GC stars - those that have been enriched by some as yet-to-be-determined process (Holtzman et al., 2015; Jönsson et al., 2018). Fernández-Trincado et al. (2020) found significant differences in certain abundance ratios compared to ASPCAP, in the range of dex higher than ASPCAP values for [N, O, Al, Si, Ce/Fe] to dex lower for [O, Mg/Fe] in 2P stars. CAPOS I showed that even the ASPCAP Fe abundances of 2P stars deviates from that of their first population (1P) cousins, with the former offset to somewhat higher metallicity values within a given GC. Hence, different abundance determination techniques are necessary for 2P stars at least, and preferably one should also perform the same analysis on 1P stars in order to ensure homogeneity.
In this paper we focus on a relatively poorly studied BGC, NGC 6316. This cluster was observed as part of CAPOS along with another BGC, NGC 6304, in the same APOGEE field, in order to maximize the number of CAPOS clusters. Another SDSS-IV program also observed stars in this same general area, and some of those stars also ended up being NGC 6316 members. Mészáros et al. (2020) notes the existence of only a single star of NGC 6316 in their large-scale study of GCs with APOGEE, but did not publish any abundances. The total number of APOGEE members studied here, 10, provides a good sample for not only deriving mean abundances but also investigating MPs.
NGC 6316 is regarded as one of the classic metal-rich BGCs, with a metallicity of 0.45 listed by Harris (1996, 2010 update). However, due to the extinction problem described above, no previous high resolution spectroscopic study has been performed on it, and it is not part of the HST UV Legacy survey, so nothing is known of its MP nature. Its metallicity has only been estimated via relatively uncertain photometric or low resolution spectroscopic means (e.g. Conroy et al. 2018, [Fe/H] = -0.87). The best of these, until now, is the Calcium triplet study of Vásquez et al. (2018), who found a mean metallicity of 0.37 0.17 from only 3 stars, in good agreement with the Harris catalog. Most recently, Deras et al. (2023) used deep HST photometry to estimate the age and metallicity, deriving [Fe/H] 0.9 and 13.1 0.5 Gyr, making it ”another extremely old relic of the assembly history of the Galaxy” according to them, but they lament the lack of a high-quality metallicity determination for this GC, which of course is critical for deriving an accurate age. Note that their value is about 0.5 dex lower than previous estimates, further demanding a more robust metallicity derivation for this cluster.
According to the Baumgardt database (and references therein), NGC 6316 lies at J2000 coordinates RA: 259.15542 deg and DEC: 28.14011 deg, it is located at a heliocentric distance of kpc and Galactocentric distance of 3.16 0.36 kpc, has mean absolute proper motions (PMs) of pmRA = 4.974 0.028 masyr-1 and pmDEC = 4.605 0.027 masyr-1, a mean heliocentric radial velocity (RV) of kmsec-1, a mass of M⊙, and a tidal radius (TR) pc. Deras et al. (2023), provides a foreground reddening of E(B-V) .
Massari, Koppelman, & Helmi (2019) classify it as a Main Bulge GC, Pérez-Villegas et al. (2020) as Thick Disk, Callingham et al. (2022) as most likely originating from Kraken, and both Belokurov & Kravtsov (2024) and Chen & Gnedin (2024) as in situ. Geisler et al. (in prep) give NGC 6316 an 81% probability of being an in situ GC and a 51% probability of being a BGC as opposed to Disk GC, and finally classify it as a bonafide BGC.
Here we present the data used, how it was obtained, reduced, and members selected (Section 2). The derivation of atmospheric parameters and the methodology used to obtain the abundances appear in section 3. The description of the abundances, analysis, comparison with the literature and analysis of its in-situ nature is given in section 4, while Section 5 contains the summary of the paper.
The final sample is depicted in all plots as bigger circles with a color code depending on their SNR.
2 Data and target selection
The data consists of high-resolution spectra (R 22500), centered in the near-infrared (H-Band), ranging from 1.51 m to 1.70 m observed through the APOGEE spectrograph (Wilson et al., 2019) on the 2.5m Irénée du Pont telescope at Las Campanas Observatory (Bowen & Vaughan, 1973), used by the APOGEE survey to observe the Southern Hemisphere (APOGEE-2S, Santana et al., 2021).
NGC 6316 was observed as part of the CAPOS survey (Geisler et al., 2021) through the Contributed APOGEE-2S CNTAC program CN2019B. Since the observations were obtained prior to the release of Gaia DR2 data, targets were based on existing position, color-magnitude and radial velocity information.
The final cluster members were selected from the final version of the APOGEE-2 catalog published in December 2021 as part of the Data Release 17 of the Sloan Digital Sky Survey (SDSS) (Abdurro’uf et al., 2022), which is available in the SDSS IV Science Archive Server222https://www.sdss4.org/dr17/irspec/spectro_data/.
Based on the information given in this catalog, we selected targets as potential members of NGC 6316 through 5 criteria:
- Using the central RA/DEC coordinates, distance and physical tidal radius TR of the cluster from Baumgardt database, we determined the angular size of the TR (0.24 deg) to select all the observed APOGEE stars inside it (Red dots in the top left panel of Figure 1)
- We took the mean PMs values from GAIA (Gaia Collaboration et al., 2016, 2021), attached to the APOGEE catalog, and selected all the stars inside a radius of 0.5 masyr-1 from the mean value given by Baumgardt database. (Blue dots in the top right panel of Figure 1)
- Taking into account the mean RV and [Fe/H] of the cluster given by Baumgardt database and Deras et al. (2023), we selected a range of stars whose RVs, determined by the APOGEE survey, were around this value, ranging from kmseg-1, and whose metallicities [Fe/H] from ASPCAP were close to the value given by Deras et al. (2023) for the cluster. (Green dots and other non blue circles in the bottom left panel of Figure 1)
- A Color-magnitude diagram with GAIA photometry in vs was made to ensure that all the potential targets lay along the main Red Giant Branch (RGB). (Colored points in bottom right panel of Figure 1)
- Finally, based on the minimum SNR suggested by APOGEE (SNR 70)333https://www.sdss4.org/dr17/irspec/abundances/, 10 out of the 19 potential targets were included in the final membership selection.
A cross-match of our 10 targets with the Vasiliev & Baumgardt (2021) catalog of NGC 6316 determined that, according to the criteria used in their study, each star has a probability of over 80% of being part of the cluster, adding further reliability that our membership selection is very robust.
Table 1 showlists the member parameters. These selections led to the previously mentioned total of 10 targets, of which 8 were observed as part of the CAPOS survey (Geisler et al., 2021), and the remaining 2 were observed under the main APOGEE survey (Majewski et al., 2017). The spectra of these 10 targets were downloaded from the SDSS IV Science Archive Server444https://data.sdss.org/sas/dr17/apogee/spectro/aspcap/dr17/synspec_rev1/. All the obtained spectra had already undergone reduction using the APOGEE-2 data reduction pipeline as described in Nidever et al. (2015).
ID | logg | [Fe/H] | [C/Fe] | [N/Fe] | [O/Fe] | [Mg/Fe] | [Al/Fe] | [Si/Fe] | ||
2M171638642809385 | 4119 | 1.188 | 2.485 | |||||||
2M171652352809502 | 4118 | 1.187 | 1.670 | |||||||
2M171636232808067 | 3842 | 0.682 | 2.950 | |||||||
2M171633302808396 | 3979 | 0.932 | 2.243 | |||||||
2M171640482808443 | 4051 | 1.062 | 2.173 | |||||||
2M171644822808302 | 4083 | 1.123 | 1.690 | |||||||
2M171633932811052 | 3884 | 0.758 | 2.699 | |||||||
2M171639032807212 | 3871 | 0.734 | 2.798 | |||||||
2M171639112804506 | 3754 | 0.522 | 2.226 | |||||||
2M171636272807166 | 3668 | 0.364 | 2.718 | |||||||
Mean | ||||||||||
0.08 | 0.25 | 0.40 | 0.18 | 0.11 | 0.14 | 0.05 | ||||
ID | [/Fe] | [P/Fe] | [K/Fe] | [Ca/Fe] | [Ti/Fe] | [V/Fe] | [Cr/Fe] | [Mn/Fe] | [Ni/Fe] | [Ce/Fe] |
2M171638642809385 | ··· | |||||||||
2M171652352809502 | ··· | ··· | ··· | |||||||
2M171636232808067 | ··· | ··· | ||||||||
2M171633302808396 | ··· | |||||||||
2M171640482808443 | ··· | |||||||||
2M171644822808302 | ||||||||||
2M171633932811052 | ··· | |||||||||
2M171639032807212 | ||||||||||
2M171639112804506 | ||||||||||
2M171636272807166 | ||||||||||
Mean | ||||||||||
0.07 | 0.09 | 0.09 | 0.10 | 0.11 | 0.11 | 0.11 | 0.10 | 0.05 | 0.15 |
3 Methodology
3.1 Atmospheric parameters
The atmospheric parameters of our targets were obtained through an interactive procedure based on the GAIA G, Gbp, Grp and 2MASS J, H, Ks photometry of the cluster. During this procedure, the effective temperature (T) and gravity (log g) of each RGB star (including our targets) were obtained, and at the same time, CMDs were corrected for differential reddening. First, we fitted a PARSEC isochrone (Bressan et al., 2012) to the RGB and Red Clump (RC), assuming an age of 13.0 Gyrs. We accounted for reddening by applying the Cardelli et al. (1989) reddening law to the isochrone. The visual absorption AV, the RV parameter, the intrinsic distance modulus (m-M)0 and the global metallicity [M/H] were determined by simultaneously fitting the RGB, RC and RGB-tip in the Ks vs. Gbp - Ks, G vs. Gbp - Grp, Gbp vs. Gbp - Grp and Ks vs. J - Ks CMDs. The T and log g of each RGB star were then determined as those corresponding to the point on the isochrone where the magnitude matches that of the star (we avoided using RC stars for this step). We used the Ks magnitude because it is the least affected by reddening and, consequently, by differential reddening. Having the temperature we obtained the intrinsic Gbp - Ks color of each star from the color-temperature relation of the RGB part of the isochrone and, by subtracting the mean reddening obtained from the isochrone fitting, also the differential reddening at the position of each star. Finally, for each star we selected the 4 closest neighbours (5 stars in total) and corrected its G, Gbp, Grp, J, H, Ks magnitudes using the mean differential reddening of the 5 stars. We used the Gbp - Ks color because it is the most sensitive to any reddening variation. This procedure was iterated until any improvement of the CMDs was negligible.
We obtained (m-M)0 = 15.32 0.05, AV = 1.91 0.05 and RV = 2.7 0.1, resulting in E(B-V) = 0.71, close to the Deras et al. (2023) value of 0.64 but substantially higher than the Harris (1996, 2010 edition) value of 0.54. This value is higher than any other from the literature since the RV we estimated (2.7) is lower than the value usually assumed (RV = 3.1). We also obtained [M/H] = 0.6 0.05, a value higher than the iron content of the cluster ([Fe/H] = 0.9) given by Deras et al. (2023) (and corroborated by us). This is not surprising since globular clusters are usually -enhanced. Also, the isochrone fitting gave an age estimation of 12 Gyrs, however, we cannot strongly rely on this estimate since no turn off point (TO) is visible in the Gaia CMDs. However, the deep HST data of Deras et al. (2023) do indeed reveal the TO and an age of 13.1 0.5 Gyr.
3.2 Chemical abundances
The method used to derive the elemental abundances was a local thermodynamics equilibrium (LTE) analysis using all the 10 targets with the Brussels Automatic Code for Characterizing High accuracy Spectra: BACCHUS (Masseron, Merle, & Hawkins, 2016)), which relies on the radiative code Turbospectrum (Alvarez & Plez, 1998; Plez, 2012) and the MARCS model atmosphere grid Gustafsson et al. (2008). With the effective temperature and surface gravity photometrically derived and an initial metallicity estimate from Deras et al. (2023) we allowed BACCHUS to determine and the convolution. The is obtained by minimizing the trend of Fe abundances against their reduced equivalent width, while the convolution parameter accounts for the total effect of instrument resolution, macroturbulence, and on the line broadening (Hawkins et al., 2016).
We then computed the elemental abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce following the same methodology as described in (Hawkins et al., 2016), briefly outlined here: First, to fit the local continuum, we performed a synthesis using the SDSS-III APOGEE spectral line list for H-band spectroscopy (Shetrone et al., 2015) updated with the Nd II and Ce II lines from Hasselquist et al. (2016) and Cunha et al. (2017), respectively. After that, we removed cosmic rays and telluric lines, and estimated the local SNR. Finally, before determining the abundances, a series of flux points contributing to a given absorption line was automatically selected.
Four different abundance determination methods are used, that are: line-profile fitting, an equivalent-width comparison, core line-intensity comparison and global goodness-of-fit estimate. For each method, we used a set of synthetic spectra of different abundances to compute the observed abundance. Only the lines with the best fit of the individual abundances are kept. We chose the goodness-of-fit diagnostic for the final abundance determination due to its reliability and consistency with our previous work. Nevertheless, we preserved data from the other diagnostic methods. To determine the C, N, and O elemental abundances, a mix of heavily CN-cycled and -rich models were used, as well as the molecular lines mentioned earlier. Oxygen abundances are estimated from the hydroxide molecular lines. Then, with these abundances in hand, we derived the carbon abundances from the carbon monoxide molecular lines, and last, the nitrogen abundances are obtained from the cyanogen molecular lines. We repeated this cycle until any dependence on OH, CO, and CN lines was negligible. After this process is finished, the rest of the mentioned elements are derived using the same four abundance determination methods.
Figure 2 shows the fit obtained for Oxygen, Aluminum and Nickel for the star 2M171639032807212 as an example of the quality of the fit described. In the Figure, the black dotted line is the observed spectrum of the target and the red continuum line is the created synthetic spectrum.
To determine and weight the errors we determined the abundances for each star again, varying only one of the parameters Teff, log g, vt and [Fe/H] by 100K, 0.3, 0.05 Kms-1 and 0.05, respectively, for a total of 8 different abundances for each element. These values were chosen as they represent the typical uncertainty in the atmospheric parameters for our sample (see e.g. Romero-Colmenares et al., 2021; González-Díaz et al., 2023). A standard deviation was calculated for each parameter separately (, , and ), including the real abundance for each element with the 2 resulting abundances for the varied parameter, in the form:
The final error was determined by taking the square root of the quadratic sum of these four different errors:
We finally took the errors from all the elements of all the stars and associated them with the corresponding abundances.
4 Results and discussion
4.1 Results and comparison with other in-situ clusters
We report abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Fe, Ni and Ce. These are shown in Table 2 together with overall [/Fe] abundances for each star determined with the available elements (O, Mg, Si and Ca), weighted equally. Each abundance has its associated weighted error, determined by the square root of the quadratic sum of each individual error and dividing the result by the number of stars. Table 2 also contains the microturbulence velocities obtained by BACCHUS. This is the first spectroscopic report of abundances manually determined for this cluster and, since we used high resolution spectroscopy, the most robust study to the date. We obtain a mean metallicity for the cluster of [Fe/H] = 0.87 0.02, with no evidence of any intrinsic Fe abundance variation. Our metallicity is in very good agreement with that obtained by Deras et al. (2023), [Fe/H] = , and Conroy et al. (2018), [Fe/H] = . However, it is some 0.5 dex lower than the canonical value of 0.45 from the Harris (1996, 2010 edition) catalog and similarly offset from the Vásquez et al. (2018) value of 0.37. Our determination involves a large sample of stars observed at high resolution and SNR, and thus should be preferred to previous estimates.
With our new metallicity, NGC 6316 is now no longer a member of the classic metal-rich BGCs. Indeed, based on both the Bica et al. (2024) and Geisler et al. (in prep) metallicity distribution functions of BGCs, which (assuming that it is indeed a BGC) are bimodal with peaks near -0.5 and -1.1, NGC 6316 falls in the lower metallicity domain, which is now known to be the dominant mode.
Given our good agreement with the Deras et al. (2023) estimate, we note that this gives increased support to their age determination of 13.1 +/- 0.5 Gyr, ensuring that NGC 6316 is indeed a rather old in situ GC for its metallicity. It now falls nicely along the age-metallicity sequence of in situ GCs, which are substantially more metal-rich than their accreted cousins of similar age, and indeed older than similar metallicity ex situ GCs (Cohen et al., 2021; Belokurov & Kravtsov, 2024). If instead the old Harris (1996, 2010 edition) metallicity was used, NGC 6316 would be a substantial outlier in both of these diagrams.
Figure 3 shows abundance relations for light elements involved in the MP phenomenon. Alongside NGC 6316 (white stars) there are four other clusters: NGC 6522 in emerald inverted triangles ([Fe/H] = , Fernández-Trincado et al. 2019), NGC 6558 in blue squares ([Fe/H] = , González-Díaz et al. 2023), Ton 2 in red triangles ([Fe/H] = , Fernández-Trincado et al. 2021) and NGC 6121 (M4) in Fuchsia circles ([Fe/H] = , Mészáros et al. 2020. These clusters were selected because all of their abundances were determined using BACCHUS, and they present similar metallicities as our analyzed stars. These criteria were selected in an attempt to find similar (or different) behaviours at similar/different location for in-situ GCs. According to Geisler et al. (in prep), all 5 of these GCs are in-situ, NGC 6316, NGC 6522 and NGC 6558 are indeed from the Bulge, Ton 2 is from the Disk, and it is uncertain whether NGC 6121 is from the Bulge or Disk.
As seen in Figure 3 top left panel, NGC 6316 presents a strong C-N anti-correlation, as do NGC 6522, NGC 6558, Ton 2 and NGC 6121, but our GC presents the strongest and most extended anticorrelation of all the compared B/D clusters. Fernández-Trincado et al. (2021) report a correlation between Ce and N in Tonantzintla 2. We observe a similar correlation, with substantial scatter, in our cluster (bottom right panel), suggesting that this cluster may have received contamination from external sources. There is a significant spread in Al and Mg, with the latter spread being higher than any of the BGC compared in the plots. This spread in both elements is expected, as Carretta et al. (2009) establishes that Al-rich and Mg-depleted stars are present in clusters that are massive (M , quite metal-poor ([Fe/H] ), or both. Our cluster, although it is not considered ”metal poor”, has a mass of according to the Baumgardt database, which is sufficient to be considered massive. However, the bottom left panel does not show a linear Mg-Al anticorrelation but rather a scatter in both elements. Mészáros et al. (2020) defined 2P as [Al/Fe] . However, this definition is not reliable for our case, since two of the three stars above this limit are the lowest in Nitrogen (hence deemed 1P) in our sample, using the definition of [N/Fe] from Geisler et al. (2021). It is worth noting that these two nitrogen-low stars are also the highest in magnesium in the sample (top right panel in Figure 2. Taken as a whole, NGC 6316 displays most of the usual symptoms of the MP disease, but this is the first time it has been diagnosed in detail.
Figure 4 shows a violin diagram comparing C, N, O, Mg, Al, Si, Ca, Fe, and Ce in NGC 6316 with those of the same clusters as in Figure 3: NGC 6522, Ton 2, NGC 6558 and NGC 6121. This diagram illustrates the range for each element, represented by a vertical line with two small horizontal lines at the limits and a third one in-between showing the mean abundance of the element. The color width of each element across the vertical line indicates the number of stars with each abundance, revealing how many stars contribute to the spread in each element.
Our cluster shows a higher spread than the other in situ GCs in most of the elements shown in the diagram, with the exception of Si. Additionally, there is generally good agreement among these GC abundances and spreads. M4 presents higher mean abundances than all the other GCs in O, Mg, Al, Si and Ce.
4.2 Comparisons with ASPCAP
Given the substantial number of new abundances obtained, it is a good opportunity to compare them with the ASPCAP values provided in the APOGEE value-added catalogue from Schiavon et al. (2024), but also taking into account that a more complete and detailed analysis is being carried out in Geisler et al. (in prep.) for all CAPOS clusters. Schiavon et al. (2024) presented abundances for a group of stars in NGC 6316, which includes all 10 of our target stars, enabling a reliable comparison. For brevity, we will refer to our results as BACCHUS and those from Schiavon et al. (2024) as ASPCAP.
The mean iron abundance in BACCHUS (0.87 0.02) is approximately 0.1 dex lower than ASPCAP (0.78 0.005), yet still aligns with the metal-poor side of the measurements when comparing with the old references. ASPCAP also presents a lower metallicity dispersion than BACCHUS. One of the most notable differences between the ASPCAP and BACCHUS abundances lies in Carbon. Although both methods show the same anticorrelation in C and N and similar trend, the two most Carbon-depleted stars, identified as 2P stars, are 0.3-0.4 dex more depleted in BACCHUS than in ASPCAP. For Nitrogen, while ASPCAP and BACCHUS exhibit similar dispersion, ASPCAP reports a mean nitrogen abundance that is 0.1 dex lower than BACCHUS. Regarding the two mentioned 2P stars, their mean nitrogen values are similar in both methods, but the BACCHUS abundances show significantly higher dispersion.
The -elements O, Mg, Si, and Ca show a difference of 0.05 in their overall [/Fe] abundance, with BACCHUS reporting a mean of [/Fe] = 0.30 0.02 (Figure 5), while ASPCAP reports a mean [/Fe] = 0.25. Notable differences arise in individual abundances. For instance, oxygen shows similar mean values between ASPCAP ([O/Fe] = 0.28) and BACCHUS ([O/Fe] = 0.26), but BACCHUS exhibits significantly higher dispersion (0.04 vs 0.18). Since MPs are present in this cluster, an oxygen dispersion is expected, lending more credibility to our individual oxygen abundances. Indeed, a C-O correlation (or N-O anticorrelation) is not observed using ASPCAP values.
Magnesium shows a strong agreement in its mean abundance between the two methods, but BACCHUS again exhibits higher dispersion. Notably, two stars in BACCHUS are more magnesium-depleted than any ASPCAP star. Among these, one also displays the high carbon depletion and enhanced nitrogen mentioned earlier, while the other presents a normal level of carbon but enhanced nitrogen abundance (and thus both considered as 2P). Aluminum and Potassium show excellent agreement in both their mean abundance and dispersion across methods. Both ASPCAP and BACCHUS report the same dispersion for Silicon, but BACCHUS abundances are 0.1 dex higher. Calcium in BACCHUS has a slightly higher abundance, with dispersion also being twice as large as in ASPCAP. Titanium shows similar dispersion in both methods but ASPCAP reports abundances that are 0.05 dex higher. However, the typical errors of our Ti measurements allow to determine an agreement between ASPCAP and BACCHUS for this element.
The iron-peak elements chromium, manganese and nickel are in excellent agreement between both methods, both in terms of abundance and dispersion. Vanadium in BACCHUS is 0.2 dex higher than in ASPCAP, though its dispersion is twice as large in the latter. Cerium shows half the dispersion and abundance in ASPCAP compared to BACCHUS, but both methods reveal the correlation with nitrogen.
It is worth noting that Vanadium is the only element that presents higher dispersion in ASPCAP abundances than in BACCHUS. 7 elements (C, N, O, Mg, Ca, Fe and Ce) present lower dispersion with the first method, while the other 7 elements (Al, Si, K, Ti, Cr, Mn and Ni) present the same dispersion with both methods. No P abundances are given by ASPCAP.
It should be noted that using our [Fe/H] and [/Fe] abundances to calculate the global metallicity [M/H] via the formula from Salaris, Chieffi, & Straniero (1993), we obtain [M/H] = -0.66 0.03. This result is very close to our photometric determination in section 3.1, further supporting the validity of our abundance and atmospheric parameter determinations. Conversely, the global metallicity reported by ASPCAP is almost identical to their iron metallicity, differing by only 0.01 dex, which does not align with their calculation using the aforementioned formula.
4.3 The in-situ nature of NGC6316
Bottom: Mg vs Al-planes but using BACCHUS abundances from Mészáros et al. (2020) and CAPOS papers. The abundances were systematically shifted as indicated in the appendix A of Belokurov & Kravtsov (2024). Left panel shows Circles representing single stars, while diamonds represents mean abundances for each cluster. Color dashed lines are the contours of the distribution of the single stars. The distribution of Mg and Al are shown in histograms at the right and top of the plot, respectively. Single star distributions show two clear dense areas where the in-situ and ex-situ stars are concentrated. This is more difficult to reproduce with mean values. The right panel shows individual stars of some CAPOS clusters compared in the previous section. The stars tend to cluster in an intermediate region between the two denser areas, slightly below the separation line in the ex-situ zone. Using only 1P stars, NGC 6316 and Ton 2 fall in the in-situ region.
As mentioned, Massari, Koppelman, & Helmi (2019) and Geisler et al. (in prep) classify NGC 6316 as an in-situ GC. Belokurov & Kravtsov (2024, hereafter BK24) proposed an abundance-based classification scheme for in-situ and ex-situ GCs, utilizing [Mg/Fe] vs [Al/Fe] abundances derived from ASPCAP data and other sources scaled to ASPCAP (see their Fig. 11, right panel). We replicated this scheme (Figure 6, top left panel) using eight GCs—four in-situ (NGC 6316, NGC 6388, NGC 6397, and Ton 2) and four ex-situ (NGC 288, NGC 1851, NGC 2808, and NGC 3201)—based on the classification given by Massari, Koppelman, & Helmi (2019), with [Mg/Fe] and [Al/Fe] abundances determined using BACCHUS (Mészáros et al. 2020 and related studies discussed earlier in this section) shifting the abundances as indicated in BK24 (see their Appendix A). Our analysis focused exclusively on 1P stars, as both Mg and Al elements are influenced by MPs. The selection of 1P stars was done trying to find an agreement with the [Al/Fe] = 0.3 limit suggested by Mészáros et al. (2020), with the lowest [N/Fe] and highest [C/Fe] stars, giving priority to the [N/Fe] and [C/Fe] levels given any contradiction with [Al/Fe]. The division line between in- and ex-situ proposed by BK24 was also included. NGC 6558 and NGC 6522 were excluded from this plot because, despite having at least one 1P star with [Al/Fe] ¡ 0.3, their [N/Fe] levels were too high. Using BACCHUS abundances, we found that the trend line proposed by BK24 did not function as expected. The separation between in-situ and ex-situ GCs was unclear, and although NGC 6316 appears in the ”in-situ” side of the plot, the separation among in- and ex-situ GCs seems to be more complicated than simply using a line.
BK24 also mentions that including 2P stars should not significantly affect the plot, as the anomalously low [Mg/Fe] abundances are counteracted by elevated [Al/Fe] levels. This shifts the GCs in the plot from the top-left to the bottom-right, parallel to the dividing line, while keeping them on the same side as when using only 1P stars. Based on this statement, we included both 1P and 2P stars from these GCs, as well as other GCs whose samples did not appear to contain 1P stars. (Figure 6, top right panel)
We found that although adding 2P stars does not drastically change the distribution of the majority of GCs in the plot, it does for NGC 6316, which exhibits MP characteristics (”MP disease”) M13 and NGC 6397, that change their positions to the other side. Thus, the addition of 2P stars is not reliable. Also, the dividing line proposed by BK24 continues to perform poorly when using 2P stars.
Another analysis was conducted using all GCs from Mészáros et al. (2020) with more than five stars, systematically shifting the data following the BK24 methodology to align the [Mg/Fe] and [Al/Fe] abundances with the BK24 separation line distinguishing in-situ and ex-situ GCs (Figure 6, bottom left panel). In there, circles represent individual stars from Mészáros et al. (2020), while diamonds indicate the mean abundance for each cluster. The color-dashed lines outline the distribution contours of individual stars. The upper and right subplots display histograms of the overall dataset’s [Mg/Fe] and [Al/Fe] distributions. The single-star distributions exhibit two distinct dense regions, corresponding to the concentrations of in-situ and ex-situ stars. However, this distinction is less clear in the mean values (diamond symbols), where the two groups overlap within a broader area.
Although 2P stars are included in this plot, they are not critical in this context. Stars with non-pristine chemical compositions (2P stars) tend to scatter across the plot without impacting the dense regions formed by in-situ and ex-situ 1P stars. In addition to the two primary dense regions, a secondary, less prominent cluster of in-situ stars is observed at approximately [Al/Fe] 0.3, [Mg/Fe] 0.3, closer to the main concentration of ex-situ stars. The BK24 separation line crosses through the two main distributions but judging by eye, it seems that the slope should be more vertical with BACCHUS abundances.
This plot largely reproduces BK24’s main result: in-situ and ex-situ stars can be distinguished in the [Mg/Fe] vs. [Al/Fe] diagram. However, the separation is not entirely clean.
The right panel displays the same contours and dashed lines as in the left panel. Star symbols represent stars from the CAPOS clusters compared in Figures 3 and 4. After applying the shift proposed by BK24, these stars tend to cluster in an intermediate region between the two densest areas, slightly below the BK24 separation line in the ex-situ zone. When considering the mean with all 1P and 2P stars, all GCs fall into the ex-situ region. However, when using only 1P stars, NGC 6316 and Tonantzintla 2 are located in the in-situ zone, while NGC 6558 and NGC 6522, whose stars all lie in the ex-situ zone, lack 1P stars in the sample, agreeing with, or at least not contradicting, the classification of Massari, Koppelman, & Helmi (2019) and Geisler et al. (in prep).
Although these results align with the existing literature, it is important to note a methodological contradiction: Our analysis works best (albeit with substantial confusion between classes), only while using 1P stars, discarding the reliability of using 2P stars, while the separation line defined by BK24 was derived from both 1P and 2P stars.
A more detailed study focusing on 1P stars in all CAPOS clusters is currently being conducted by Geisler et al. (in prep.).
5 Conclusions
In this paper, we have derived detailed abundances for a wide variety of elements for 10 RGB stars in the BGC NGC 6316 using the Brussels Automatic Code for Characterizing High accuracy Spectra (BACCHUS) based on CAPOS spectra. These stars were selected based on their strong affinity with the cluster’s mean spatial, orbital, and chemical parameters (TR, PMs, RVs, and ASPCAP [Fe/H]). Based on our findings, we can state the following:
1.
We obtained a mean metallicity of [Fe/H]= 0.87 0.02 for the cluster, with no clear evidence of any intrinsic Fe abundance variation, in very good agreement with the results of Deras et al. (2023). This value is 0.5 dex more metal poor than the values traditionally derived for this cluster (Vásquez et al., 2018; Harris, 1996, 2010 edition). We encourage the use of this new derived metallicity above the previous ones since our determination involves a large sample of stars observed at high resolution and SNR. Also, since our metallicity supports that derived by Deras, the age derived by them (13.1 ± 0.5 Gyrs) is therefore supported by this study.
2.
We provide abundances for C, N, O, Mg, Al, Si, P, K, Ca, Ti, V, Cr, Mn, Ni and Ce. This is the first non-automatic spectroscopic abundance determination for this cluster, hence the most reliable chemical study to date for this cluster, preferable to previous studies like Schiavon et al. (2024) and Conroy et al. (2018), who used ASPCAP and stellar population models, respectively. With the abundances of our available -elements we derived [/Fe] = 0.3 0.02.
3.
The C and N levels of the stars show a clear anticorrelation, with a noticeable enhancement in Nitrogen, indicating the presence of MPs within the cluster.
4.
The isochrone fitting included in our method for determining the atmospheric parameters provided the following information for NGC 6316: AV = 1.91 0.05, RV = 2.7 0.1 (hence giving a higher E(B-V) = 0.71 than other studies that use RV = 3.1), (m-M)0 = 15.32 0.05 and [M/H] = 0.6 0.05. This latter is in good agreement with our global metallicity derived from our [Fe/H] and [/Fe] abundances ([M/H] = 0.66 0.03) using the formula from Salaris, Chieffi, & Straniero (1993).
5.
A comparison of the abundances of NGC 6316 with those of other in-situ GCs shows a very good agreement in almost all cases, with NGC 6316 having higher spread in most cases except Al and Si.
6.
Although it is known that ASPCAP does not perform well for 2P stars, inflating the precision of mean abundances, the elements Al and K, and the iron-peak elements Ti, Cr, Mn and Ni show excellent agreement in the mean abundances and dispersion for our GC, while O and Mg show a good agreement only in their mean abundances. Vanadium is the only element that presents higher dispersion in ASPCAP abundances than in BACCHUS, while all the other elements present lower (C, N, O, Mg, Ca, Fe and Ce) or equal (Al, Si, K, Ti, Cr, Mn and Ni) dispersions than BACCHUS abundances.
7.
Replicating the scheme proposed by Belokurov & Kravtsov (2024), but based on the mean [Mg/Fe] and [Al/Fe] abundances of 1P stars only, NGC 6316 is classified as an in-situ GC. This study discards the reliability of using 2P stars. However, the separation line defined by BK24 was derived from both 1P and 2P stars. The authors acknowledge that a separation line derived exclusively from 1P stars should be implemented to ensure a more consistent and accurate analysis.
Acknowledgements.
SV gratefully acknowledges the support provided by Fondecyt regular no. 1220264 and by the ANID BASAL projects FB210003. D.G. gratefully acknowledges the support provided by Fondecyt regular n. 1220264. D.G. also acknowledges financial support from the Dirección de Investigación y Desarrollo de la Universidad de La Serena through the Programa de Incentivo a la Investigación de Académicos (PIA-DIDULS). The team acknowledges J.G. Fernandez-Trincado for his help and commentaries. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta K., Aerts C., Silva Aguirre V., Ahumada R., Ajgaonkar N., Filiz Ak N., et al., 2022, ApJS, 259, 35. doi:10.3847/1538-4365/ac4414
- Allende Prieto et al. (2008) Allende Prieto C., Majewski S. R., Schiavon R., Cunha K., Frinchaboy P., Holtzman J., Johnston K., et al., 2008, AN, 329, 1018. doi:10.1002/asna.200811080
- Alonso et al. (1999) Alonso, A., Arribas, S., & Martínez-Roger, C. 1999, A&AS, 140, 261
- Alvarez & Plez (1998) Alvarez R., Plez B., 1998, A&A, 330, 1109. doi:10.48550/arXiv.astro-ph/9710157
- Bajkova & Bobylev (2020) Bajkova A. T., Bobylev V. V., 2020, arXiv, arXiv:2008.13624
- Baumgardt et al. (2019) Baumgardt, H., Hilker, M., Sollima, A., et al. 2019, MNRAS, 482, 5138. doi:10.1093/mnras/sty2997
- Bastian et al. (2013) Bastian, N., Lamers, H. J. G. L. M., de Mink, S. E., et al. 2013, MNRAS, 436, 2398
- Bastian & Lardo (2018) Bastian N., Lardo C., 2018, ARA&A, 56, 83. doi:10.1146/annurev-astro-081817-051839
- Bedin et al. (2004) Bedin, L. R., Piotto, G., Anderson, J., et al. 2004, ApJ, 605, L125
- Bellazzini, Ferraro, & Ibata (2003) Bellazzini M., Ferraro F. R., Ibata R., 2003, AJ, 125, 188. doi:10.1086/344072
- Belokurov & Kravtsov (2024) Belokurov V., Kravtsov A., 2024, MNRAS, 528, 3198. doi:10.1093/mnras/stad3920
- Bessell (2005) Bessell M. S., 2005, ARA&A, 43, 293. doi:10.1146/annurev.astro.41.082801.100251
- Bica et al. (2024) Bica E., Ortolani S., Barbuy B., Oliveira R. A. P., 2024, A&A, 687, A201. doi:10.1051/0004-6361/202346377
- Bonatto, Campos, & Kepler (2013) Bonatto C., Campos F., Kepler S. O., 2013, MNRAS, 435, 263. doi:10.1093/mnras/stt1304
- Bovy et al. (2019) Bovy, J., Leung, H. W., Hunt, J. A. S., et al. 2019, MNRAS, 490, 4740. doi:10.1093/mnras/stz2891
- Bowen & Vaughan (1973) Bowen I. S., Vaughan A. H., 1973, ApOpt, 12, 1430. doi:10.1364/AO.12.001430
- Bressan et al. (2012) Bressan A., Marigo P., Girardi L., Salasnich B., Dal Cero C., Rubele S., Nanni A., 2012, MNRAS, 427, 127. doi:10.1111/j.1365-2966.2012.21948.x
- Callingham et al. (2022) Callingham T. M., Cautun M., Deason A. J., Frenk C. S., Grand R. J. J., Marinacci F., 2022, MNRAS, 513, 4107. doi:10.1093/mnras/stac1145
- Campbell et al. (2012) Campbell S. W., Yong D., Wylie-de Boer E. C., Stancliffe R. J., Lattanzio J. C., Angelou G. C., D’Orazi V., et al., 2012, ApJL, 761, L2. doi:10.1088/2041-8205/761/1/L2
- Cannon & Stobie (1973) Cannon R. D., Stobie R. S., 1973, MNRAS, 162, 207. doi:10.1093/mnras/162.3.207
- Canterna (1976) Canterna, R. 1976, AJ, 81, 228
- Cardelli, Clayton, & Mathis (1989) Cardelli J. A., Clayton G. C., Mathis J. S., 1989, ApJ, 345, 245. doi:10.1086/167900
- Carretta et al. (2009) Carretta E., Bragaglia A., Gratton R. G., Lucatello S., Catanzaro G., Leone F., Bellazzini M., et al., 2009, A&A, 505, 117. doi:10.1051/0004-6361/200912096
- Carretta et al. (2009) Carretta E., Bragaglia A., Gratton R., Lucatello S., 2009, A&A, 505, 139. doi:10.1051/0004-6361/200912097
- Carretta et al. (2010) Carretta E., Bragaglia A., Gratton R. G., Recio-Blanco A., Lucatello S., D’Orazi V., Cassisi S., 2010, A&A, 516, A55. doi:10.1051/0004-6361/200913451
- Chen & Gnedin (2024) Chen Y., Gnedin O. Y., 2024, MNRAS, 527, 3692. doi:10.1093/mnras/stad3345
- Cohen (2004) Cohen J. G., 2004, AJ, 127, 1545. doi:10.1086/382104
- Cohen et al. (2021) Cohen R. E., Bellini A., Casagrande L., Brown T. M., Correnti M., Kalirai J. S., 2021, AJ, 162, 228. doi:10.3847/1538-3881/ac281f
- Conroy et al. (2018) Conroy C., Villaume A., van Dokkum P. G., Lind K., 2018, ApJ, 854, 139. doi:10.3847/1538-4357/aaab49
- Cummings et al. (2014) Cummings J. D., Geisler D., Villanova S., Carraro G., 2014, AJ, 148, 27. doi:10.1088/0004-6256/148/2/27
- Cunha et al. (2017) Cunha K., Smith V. V., Hasselquist S., Souto D., Shetrone M. D., Allende Prieto C., Bizyaev D., et al., 2017, ApJ, 844, 145. doi:10.3847/1538-4357/aa7beb
- D’Ercole et al. (2008) D’Ercole, A., Vesperini, E., D’Antona, F., McMillan, S. L. W., & Recchi, S. 2008, MNRAS, 391, 825
- Dalessandro et al. (2018) Dalessandro E., Lardo C., Cadelano M., Saracino S., Bastian N., Mucciarelli A., Salaris M., et al., 2018, A&A, 618, A131. doi:10.1051/0004-6361/201833650
- Decressin et al. (2007) Decressin, T., Charbonnel, C., & Meynet, G. 2007, A&A, 475, 859
- Deras et al. (2023) Deras D., Cadelano M., Ferraro F. R., Lanzoni B., Pallanca C., 2023, ApJ, 942, 104. doi:10.3847/1538-4357/aca9ce
- Dotter et al. (2011) Dotter, A., Sarajedini, A., & Anderson, J. 2011, ApJ, 738, 74
- Dotter et al. (2018) Dotter, A., Milone, A. P., Conroy, C., et al. 2018, ApJ, 865, L10
- Elmegreen (2017) Elmegreen B. G., 2017, ApJ, 836, 80. doi:10.3847/1538-4357/836/1/80
- Fernie (1983) Fernie, J. D. 1983, PASP, 95, 782
- Fernández-Trincado et al. (2019) Fernández-Trincado J. G., Zamora O., Souto D., Cohen R. E., Dell’Agli F., García-Hernández D. A., Masseron T., et al., 2019, A&A, 627, A178. doi:10.1051/0004-6361/201834391
- Fernández-Trincado et al. (2020) Fernández-Trincado J. G., Beers T. C., Minniti D., Tang B., Villanova S., Geisler D., Pérez-Villegas A., et al., 2020, A&A, 643, L4. doi:10.1051/0004-6361/202039207
- Fernández-Trincado et al. (2021) Fernández-Trincado J. G., Beers T. C., Barbuy B., Mészáros S., Minniti D., Smith V. V., Cunha K., et al., 2021, ApJL, 918, L9. doi:10.3847/2041-8213/ac1c7e
- Fernández-Trincado et al. (2022) Fernández-Trincado J. G., Villanova S., Geisler D., Barbuy B., Minniti D., Beers T. C., Mészáros S., et al., 2022, A&A, 658, A116. doi:10.1051/0004-6361/202141742
- Frelijj et al. (2017) Frelijj H., Geisler D., Cummings J., Cohen R. E., Mauro F., Munoz C., Villanova S., et al., 2017, MNRAS, 472, 4532. doi:10.1093/mnras/stx2280
- Frelijj et al. (2021) Frelijj H., Villanova S., Muñoz C., Fernández-Trincado J. G., 2021, MNRAS, 503, 867. doi:10.1093/mnras/stab443
- Frelijj et al. (2022) Frelijj H., Geisler D., Villanova S., Munoz C., 2022, RMxAA, 58, 181. doi:10.22201/ia.01851101p.2022.58.02.02
- Gaia Collaboration et al. (2016) Gaia Collaboration, Prusti T., de Bruijne J. H. J., Brown A. G. A., Vallenari A., Babusiaux C., Bailer-Jones C. A. L., et al., 2016, A&A, 595, A1. doi:10.1051/0004-6361/201629272
- Gaia Collaboration et al. (2021) Gaia Collaboration, Brown A. G. A., Vallenari A., Prusti T., de Bruijne J. H. J., Babusiaux C., Biermann M., et al., 2021, A&A, 649, A1. doi:10.1051/0004-6361/202039657
- García Pérez et al. (2016) García Pérez A. E., Allende Prieto C., Holtzman J. A., Shetrone M., Mészáros S., Bizyaev D., Carrera R., et al., 2016, AJ, 151, 144. doi:10.3847/0004-6256/151/6/144
- Geisler, Claria, & Minniti (1991) Geisler D., Claria J. J., Minniti D., 1991, AJ, 102, 1836. doi:10.1086/116008
- Geisler (1996) Geisler D., 1996, AJ, 111, 480. doi:10.1086/117799
- Geisler et al. (2021) Geisler D., Villanova S., O’Connell J. E., Cohen R. E., Moni Bidin C., Fernández-Trincado J. G., Muñoz C., et al., 2021, A&A, 652, A157. doi:10.1051/0004-6361/202140436
- Gieles et al. (2018) Gieles M., Charbonnel C., Krause M. G. H., Hénault-Brunet V., Agertz O., Lamers H. J. G. L. M., Bastian N., et al., 2018, MNRAS, 478, 2461. doi:10.1093/mnras/sty1059
- González-Díaz et al. (2023) González-Díaz D., Fernández-Trincado J. G., Villanova S., Geisler D., Barbuy B., Minniti D., Beers T. C., et al., 2023, MNRAS, 526, 6274. doi:10.1093/mnras/stad3178
- Gratton, Carretta, & Castelli (1996) Gratton R. G., Carretta E., Castelli F., 1996, A&A, 314, 191. doi:10.48550/arXiv.astro-ph/9603011
- Gustafsson et al. (2008) Gustafsson B., Edvardsson B., Eriksson K., Jørgensen U. G., Nordlund Å., Plez B., 2008, A&A, 486, 951. doi:10.1051/0004-6361:200809724
- Hamuy et al. (2006) Hamuy M., Folatelli G., Morrell N. I., Phillips M. M., Suntzeff N. B., Persson S. E., Roth M., et al., 2006, PASP, 118, 2. doi:10.1086/500228
- Han et al. (2009) Han S.-I., Lee Y.-W., Joo S.-J., Sohn S. T., Yoon S.-J., Kim H.-S., Lee J.-W., 2009, ApJL, 707, L190. doi:10.1088/0004-637X/707/2/L190
- Harris (1996) Harris W. E., 1996, AJ, 112, 1487. doi:10.1086/118116
- Hasselquist et al. (2016) Hasselquist S., Shetrone M., Cunha K., Smith V. V., Holtzman J., Lawler J. E., Allende Prieto C., et al., 2016, ApJ, 833, 81. doi:10.3847/1538-4357/833/1/81
- Hawkins et al. (2016) Hawkins K., Masseron T., Jofré P., Gilmore G., Elsworth Y., Hekker S., 2016, A&A, 594, A43. doi:10.1051/0004-6361/201628812
- Holtzman et al. (2015) Holtzman J. A., Shetrone M., Johnson J. A., Allende Prieto C., Anders F., Andrews B., Beers T. C., et al., 2015, AJ, 150, 148. doi:10.1088/0004-6256/150/5/148
- Johnson et al. (2017) Johnson C. I., Caldwell N., Rich R. M., Mateo M., Bailey J. I., Clarkson W. I., Olszewski E. W., et al., 2017, ApJ, 836, 168. doi:10.3847/1538-4357/836/2/168.
- Jönsson et al. (2018) Jönsson H., Allende Prieto C., Holtzman J. A., Feuillet D. K., Hawkins K., Cunha K., Mészáros S., et al., 2018, AJ, 156, 126. doi:10.3847/1538-3881/aad4f5
- Koch & Caffau (2011) Koch A., Caffau E., 2011, A&A, 534, A52. doi:10.1051/0004-6361/201117678
- Kurucz (1970) Kurucz, R. L. 1970, SAO Special Report, 309
- Law & Majewski (2010) Law D. R., Majewski S. R., 2010, ApJ, 718, 1128. doi:10.1088/0004-637X/718/2/1128
- Majewski et al. (2017) Majewski S. R., Schiavon R. P., Frinchaboy P. M., Allende Prieto C., Barkhouser R., Bizyaev D., Blank B., et al., 2017, AJ, 154, 94. doi:10.3847/1538-3881/aa784d
- Massari, Koppelman, & Helmi (2019) Massari D., Koppelman H. H., Helmi A., 2019, A&A, 630, L4. doi:10.1051/0004-6361/201936135
- Masseron, Merle, & Hawkins (2016) Masseron T., Merle T., Hawkins K., 2016, ascl.soft. ascl:1605.004
- Mészáros et al. (2020) Mészáros S., Masseron T., García-Hernández D. A., Allende Prieto C., Beers T. C., Bizyaev D., Chojnowski D., et al., 2020, MNRAS, 492, 1641. doi:10.1093/mnras/stz3496
- Milone et al. (2009) Milone A. P., Stetson P. B., Piotto G., Bedin L. R., Anderson J., Cassisi S., Salaris M., 2009, A&A, 503, 755. doi:10.1051/0004-6361/200912256
- Milone et al. (2017) Milone A. P., Piotto G., Renzini A., Marino A. F., Bedin L. R., Vesperini E., D’Antona F., et al., 2017, MNRAS, 464, 3636. doi:10.1093/mnras/stw2531
- Minniti et al. (2010) Minniti D., Lucas P. W., Emerson J. P., Saito R. K., Hempel M., Pietrukowicz P., Ahumada A. V., et al., 2010, NewA, 15, 433. doi:10.1016/j.newast.2009.12.002
- Monaco et al. (2018) Monaco L., Villanova S., Carraro G., Mucciarelli A., Moni Bidin C., 2018, A&A, 616, A181. doi:10.1051/0004-6361/201832979
- Moreno, Pichardo, & Schuster (2015) Moreno E., Pichardo B., Schuster W. J., 2015, MNRAS, 451, 705. doi:10.1093/mnras/stv962
- Muñoz et al. (2020) Muñoz C., Villanova S., Geisler D., Cortés C. C., Moni Bidin C., Cohen R. E., Saviane I., et al., 2020, MNRAS, 492, 3742. doi:10.1093/mnras/stz3586
- Muñoz et al. (2021) Muñoz C., Geisler D., Villanova S., Sarajedini A., Frelijj H., Vargas C., Monaco L., et al., 2021, MNRAS, 506, 4676. doi:10.1093/mnras/stab1891
- Nidever et al. (2015) Nidever D. L., Holtzman J. A., Allende Prieto C., Beland S., Bender C., Bizyaev D., Burton A., et al., 2015, AJ, 150, 173. doi:10.1088/0004-6256/150/6/173
- O’Donnell (1994) O’Donnell J. E., 1994, ApJ, 422, 158. doi:10.1086/173713
- Parmentier & Pasquali (2022) Parmentier G., Pasquali A., 2022, ApJ, 924, 81. doi:10.3847/1538-4357/ac32d8
- Pérez-Villegas et al. (2020) Pérez-Villegas A., Barbuy B., Kerber L. O., Ortolani S., Souza S. O., Bica E., 2020, MNRAS, 491, 3251. doi:10.1093/mnras/stz3162
- Piotto et al. (2015) Piotto, G., Milone, A. P., Bedin, L. R., et al. 2015, AJ, 149, 91
- Plez (2012) Plez B., 2012, ascl.soft. ascl:1205.004
- Ramírez & Meléndez (2005) Ramírez, I. & Meléndez, J. 2005, ApJ, 626, 465
- Renzini et al. (2015) Renzini A., D’Antona F., Cassisi S., King I. R., Milone A. P., Ventura P., Anderson J., et al., 2015, MNRAS, 454, 4197. doi:10.1093/mnras/stv2268
- Robin et al. (2003) Robin A. C., Reylé C., Derrière S., Picaud S., 2003, A&A, 409, 523. doi:10.1051/0004-6361:20031117
- Robin et al. (2014) Robin A. C., Reylé C., Fliri J., Czekaj M., Robert C. P., Martins A. M. M., 2014, A&A, 569, A13. doi:10.1051/0004-6361/201423415
- Romero-Colmenares et al. (2021) Romero-Colmenares M., Fernández-Trincado J. G., Geisler D., Souza S. O., Villanova S., Longa-Peña P., Minniti D., et al., 2021, A&A, 652, A158. doi:10.1051/0004-6361/202141294
- Saito et al. (2024) Saito R. K., Hempel M., Alonso-García J., Lucas P. W., Minniti D., Alonso S., Baravalle L., et al., 2024, arXiv, arXiv:2406.16646. doi:10.48550/arXiv.2406.16646
- Salaris, Chieffi, & Straniero (1993) Salaris M., Chieffi A., Straniero O., 1993, ApJ, 414, 580. doi:10.1086/173105
- Salinas, & Strader (2015) Salinas, R., & Strader, J. 2015, ApJ, 809, 169
- Sandage (1953) Sandage A. R., 1953, AJ, 58, 61. doi:10.1086/106822
- Sanders, Smith, & Evans (2019) Sanders J. L., Smith L., Evans N. W., 2019, MNRAS, 488, 4552. doi:10.1093/mnras/stz1827
- Santana et al. (2021) Santana F. A., Beaton R. L., Covey K. R., O’Connell J. E., Longa-Peña P., Cohen R., Fernández-Trincado J. G., et al., 2021, AJ, 162, 303. doi:10.3847/1538-3881/ac2cbc
- Sbordone et al. (2005) Sbordone L., Bonifacio P., Marconi G., Buonanno R., Zaggia S., 2005, A&A, 437, 905. doi:10.1051/0004-6361:20042315
- Sbordone et al. (2007) Sbordone L., Bonifacio P., Buonanno R., Marconi G., Monaco L., Zaggia S., 2007, A&A, 465, 815. doi:10.1051/0004-6361:20066385
- Sbordone et al. (2011) Sbordone, L., Salaris, M., Weiss, A., & Cassisi, S. 2011, A&A, 534, A9
- Schiavon et al. (2017) Schiavon R. P., Johnson J. A., Frinchaboy P. M., Zasowski G., Mészáros S., García-Hernández D. A., Cohen R. E., et al., 2017, MNRAS, 466, 1010. doi:10.1093/mnras/stw3093
- Schiavon et al. (2024) Schiavon R. P., Phillips S. G., Myers N., Horta D., Minniti D., Allende Prieto C., Anguiano B., et al., 2024, MNRAS, 528, 1393. doi:10.1093/mnras/stad3020
- Shetrone et al. (2015) Shetrone M., Bizyaev D., Lawler J. E., Allende Prieto C., Johnson J. A., Smith V. V., Cunha K., et al., 2015, ApJS, 221, 24. doi:10.1088/0067-0049/221/2/24
- Sirianni et al. (2005) Sirianni, M., Jee, M. J., Benítez, N., et al. 2005, PASP, 117, 1049. doi:10.1086/444553
- Smith et al. (2021) Smith V. V., Bizyaev D., Cunha K., Shetrone M. D., Souto D., Allende Prieto C., Masseron T., et al., 2021, AJ, 161, 254. doi:10.3847/1538-3881/abefdc
- Smith et al. (2002) Smith, G. H., Sneden, C., & Kraft, R. P. 2002, AJ, 123, 1502. doi:10.1086/338855
- Sneden (1973) Sneden, C. 1973, ApJ, 184, 839. doi:10.1086/152374
- Stetson (1987) Stetson P. B., 1987, PASP, 99, 191. doi:10.1086/131977
- Stetson (1994) Stetson P. B., 1994, PASP, 106, 250. doi:10.1086/133378
- Taylor (2005) Taylor, M. B. 2005, Astronomical Data Analysis Software and Systems XIV, 347, 29
- Tautvaišienė et al. (2004) Tautvaišienė, G., Wallerstein, G., Geisler, D., et al. 2004, AJ, 127, 373
- Vasiliev & Baumgardt (2021) Vasiliev E., Baumgardt H., 2021, yCat, 750. J/MNRAS/505/5978
- Vásquez et al. (2018) Vásquez S., Saviane I., Held E. V., Da Costa G. S., Dias B., Gullieuszik M., Barbuy B., et al., 2018, A&A, 619, A13. doi:10.1051/0004-6361/201833525
- Villanova et al. (2013) Villanova, S., Geisler, D., Carraro, G., Moni Bidin, C., & Muñoz, C. 2013, ApJ, 778, 186
- Walker et al. (2011) Walker A. R., Kunder A. M., Andreuzzi G., di Cecco A., Stetson P. B., Monelli M., Cassisi S., et al., 2011, MNRAS, 415, 643. doi:10.1111/j.1365-2966.2011.18736.x
- Wilson et al. (2019) Wilson J. C., Hearty F. R., Skrutskie M. F., Majewski S. R., Holtzman J. A., Eisenstein D., Gunn J., et al., 2019, PASP, 131, 055001. doi:10.1088/1538-3873/ab0075